The solubility product (Ksp) is 0.005 for the reaction below,

Ba (OH)2 (s) ⇌ Ba2+ (aq) + 2OH-(aq)


If the initial concentration of Ba (OH)2 is [Ba(OH)2] = 2M, what is the concentration of

[H3O+] at equilibrium?


1
Expert's answer
2021-06-07T03:16:05-0400

According to the equlibrium equation:

Ksp = [Ba2+]eq[OH-]2eq

Let assume that [Ba2+] = x. Then, [OH-] = 2x. Finally:

Ksp = x(2x)2

Ksp = 4x3

x3 = Ksp / 4

From here:

x = (Ksp / 4)1/3 = (0.005 / 4)1/3 = 0.107

As a result, [OH-] = 2 × 0.107 = 0.214

Finally:

[H3O+] = 1 × 10-14 / [OH-] = 1 × 10-14 / 0.214 = 4.7 × 10-14 M


Answer: 4.7 × 10-14 M